Together, research vessels A and B collected 500 seawater samples. If vessel B collected 56 more samples than vessel A, how many water samples did each vessel collect?
Answer by Theo(13342)

A + B = 500
B = A + 56
replace B with A + 56 in the first equation to get:
A + A + 56 = 500
combine like terms and subtract 56 from both sides of the equation to get:
2A = 444
solve for A to get A = 222
B = A + 56 gets you B = 278
A + B = 222 + 278 = 500, confirming the values of A and B are correct.
in these equations, A represents the number of seawater samples that vessel A collects and B represents the number of seawater samples that vessel B collects.